# DNABERT-S: Learning Species-Aware DNA Embedding with Genome Foundation Models

### v1

@@ 3.2 Curriculum Contrastive Learning ($\text{C}^{2}$LR)
[…]
@@
[…]
@@ 3.3 Implementation
In the $\text{C}^{2}$LR method, we set temperature $\tau$ as $0.05$ and hyperparameter $\alpha$ as $1.0$. We train the model for one epoch in phase I using loss Eq. (2) and for two epochs in phase II using loss Eq. (3). We use mean pooling of the last hidden states of all the tokens as the DNA embedding. We employ the Adam optimizer [Kingma and Ba, 2014], with a learning rate of 3e63e-6 and batch size of $48$. We save the model every 10000 training steps and select the best one based on the validation loss in the validation dataset. We use the pre-trained DNABERT-2 [Zhou et al., 2023] as the starting point of contrastive training. We also conduct parallel experiments with HyenaDNA [Nguyen et al., 2023]. In Appendix C.1, we show that DNABERT-2 outperforms HyenaDNA after the same contrastive training. The entire training process of DNABERT-S takes approximately 48 hours on 8 NVIDIA A100 80GB GPUs.
@@ 4 Data
4.1 Training
Each training sample of DNABERT-S is a pair of non-overlapping DNA sequences extracted from the same genome. The dataset is constructed with the reference genomes from GenBank [Benson et al., 2012]. We obtained $47923$ pairs from $17636$ viral genomes, $1$ million pairs from $5011$ fungi genomes, and $1$ million pairs from $6402$ bacteria genomes. We randomly selected $2$ million pairs from the entire $2047923$ pairs of DNA sequences to construct the training data. The rest pairs are treated as validation data. All the DNA sequences are 10000 bp in length.
4.2 Evaluation
Our evaluation spans on the Critical Assessment of Metagenome Interpretation (CAMI) II [Meyer et al., 2022] challenge benchmark and $4$ synthetic datasets. CAMI2 is one of the most comprehensive and rigorous benchmarks for metagenomics research. The datasets in CAMI2 are designed to mimic realistic microbiome environments and include a vast array of both new and known genomes, as well as plasmids and viruses. It aligns our study with real-world ecological and biological scenarios, providing a robust and contextually relevant evaluation for the DNA embedding models. We utilize $7$ datasets of long-read contigs respectively from the Marine and Plant-associated environments, where each dataset consists of 150150k-200200k DNA sequences belonging to about 100750100-750 different species sampled from $1680$ microbial genomes and $599$ circular elements. We also create $4$ Synthetic datasets by randomly extracting DNA sequences from fungi and viral reference genomes that do not overlap with our training data. Table 3 shows the statistics of the datasets we used for evaluation.
To comprehensively understand the impact of DNA embedding in various scenarios, we conduct three types of tasks: 1) species clustering given the number of species, 2) species classification with different numbers of labeled samples, and 3) metagenomics binning with an unknown number of species. The first two tasks evaluate the models in a more standard yet unrealistic setting, while the third task is a realistic problem. Since the CAMI2 datasets are highly imbalanced, for the clustering and classification tasks, we filtered the datasets to eliminate species with fewer than $100$ sequences and only kept $100$ sequences for each species, resulting in a set of perfectly balanced datasets. For the metagenomics binning problem, to mimic real-world scenarios, where the number of existing samples is unknown, we do not balance the data. Instead, following Kang et al. [2015], we only keep DNA sequences longer than 25002500bp and filter out species with fewer than $10$ sequences.
@@ 5 Experiments
In this section, we present experimental design and empirical results. We introduce baselines in Sec. 5.1 and respectively present the results of clustering in Sec. 5.2, classification in Sec. 5.3, and metagenomics binning in Sec. 5.4. In Sec. 5.5, we present ablation studies on curriculum learning and the proposed Manifold Instance Mixup training objective. We also provide empirical analysis on varying backbone models, different input lengths, reduced feature dimensions, scenarios with abundant training data, and various other types of tasks. Due to space limitations, we present them in Appendix C. For all tasks, we perform $5$ independent runs with different random seeds for each model and report the averaged results.
@@ 5.1 Baselines
We compare our model with four lines of work to examine its effectiveness in generating DNA embedding.
TNF, TNF-K, and TNF-VAE are the most widely used DNA embedding methods in metagenomics binning tools [Kang et al., 2015, 2019; Nissen et al., 2021]. TNF represents Tetra-Nucleotide Frequency, which uses the appearance frequency of each unique 4-mer ($4^{4}=256$ in total) in a DNA sequence as its embedding. TNF-K [Nissen et al., 2021] reduces TNF to $103$-dimension with a linear kernel, which utilizes DNA characteristics to reduce the correlations among different dimensions of the original TNF feature. TNF-VAE trains a Variational Autoencoder [Kingma and Welling, 2013] using TNF as input to extract features.
DNA2Vec [Ng, 2017] learns pre-trained K-mer embedding inspired by Mikolov et al. [2013]. We set $K=4$ to make it directly comparable with TNF and use the average of the $4$-mer embeddings as the DNA embedding.
DNABERT-2 [Zhou et al., 2023], HyenaDNA [Nguyen et al., 2023], and NT (Nucleotide Transformer) [Dalla-Torre et al., 2023] are representative genome foundation models. We use the average of the last hidden states as the DNA embedding. For evaluations, we utilized the respective pre-trained models from Huggingface ModelHub, specifically zhihan1996/DNABERT-2-117M, LongSafari/hyenadna-medium-450k-seqlen-hf, and InstaDeepAI/nucleotide-transformer-v2-100m-multi-species.
DNA-Mutate, DNA-Dropout, and DNA-Double are variants of DNABERT-S, with the same hyperparameters and starting checkpoint but different positive pair construction strategies in contrastive training. DNA-Mutate views the same DNA sequence before and after random mutation (i.e., swap and delete $5\%$ of nucleotides) as a positive pair. DNA-Dropout is inspired by Gao et al. [2021], which passes the same DNA sequence through the embedding model (with a dropout rate as $0.1$) twice and views the two distinct embeddings as a positive pair. DNA-Double views a DNA sequence and its complementary (e.g., AATTC v.s. TTAAG) as a positive pair.
@@ 5.2 Clustering
In this task, we evaluate the embedding quality by how well a standard clustering algorithm can distinguish and cluster different species based on the embedding. To reduce the effects of other factors, we assume the number of species is known in this task. For each dataset, we compute the embedding of each DNA sequence and perform K-means clustering by setting the num_clusters as the number of species that exist in this dataset. We employ the Adjusted Rand Index (ARI) as the evaluation metric. ARI is a measure of the similarity between two data clusterings, adjusted for chance, providing a normalized index that ranges from 1-1 to $1$. A higher ARI score indicates better performance.
Table 1 shows the models’ performance on clustering. As shown in the table, DNABERT-S consistently achieves the best performance on all the datasets and doubles the performance of the strongest baseline on average. Among all the baselines, TNF and its variant TNF-K achieve the best performance, explaining their wide usage in metagenomics binning. Yet, TNF’s performance is heavily limited since it is not learnable. TNF-VAE represents a naive algorithm that enables learning with TNF, yet it leads to big performance degradation, potentially resulting from the large gap between its training objective and the specific downstream application. Similarly, pre-trained Kmer embeddings from DNA2Vec also fail to effectively cluster different species.
Existing genome foundation models training with language modeling objectives, such as HyenaDNA and DNABERT-2, despite their remarkable performance on labeled datasets, also fail to generate representative embedding without fine-tuning. The phenomenon that pre-trained foundation models underperform descriptive textual features in generating embedding for clustering and retrieval is also observed in the field of natural language processing [Reimers and Gurevych, 2019].
@@ 5.3 Classification
In this task, we evaluate the embedding quality by how well a linear model can classify different species based on a few labeled embeddings. This evaluation is also known as linear probing. As shown in Table 3, all the datasets we use for classification consist of $100$ DNA sequence for each species. We first compute the embedding of each DNA sequence with each model. In each evaluation run, we independently select $80$ embeddings from each species to form the test set. For the rest DNA sequences, we respectively sample $1$, 22, 55, $10$, and $20$ embeddings from each species to form the training set. A Logistic Regression model is trained on the training set and evaluated on the test set. We use the macro F1 score as the evaluation metric.
Figure 3 shows the models’ performance on $6$ datasets. The remaining results are consistent and are presented in Appendix C.2. As shown in the figure, DNABERT-S consistently achieves the best performance. Remarkably, DNABERT-S achieves better performance than the strongest baseline with only $20\%$ of training data. For example, with only $2$ training samples per category, DNABERT-S achieves higher F1 scores than the strongest baseline with $10$ training samples. With the same amount of training samples, DNABERT-S significantly outperforms the baselines by a large gap. Notably, in the Synthetic datasets, where none of the species are seen during the contrastive training, a linear model trained with DNABERT-S embeddings achieves an F1 score of over $0.8$ in 200 classes classification with only $5$ labeled samples in each species, showing DNABERT-S’s capability in generalizing well on unseen data.
@@
5.4 Metagenomics Binning
Metagenomics binning is a crucial process in microbial ecology, involving the categorization of DNA sequences into groups that represent individual species. State-of-the-art metagenomics binning method [Kang et al., 2015, 2019; Nissen et al., 2021] always formulate this problem as a clustering problem with an unknown number of clusters based on the feature of each DNA sequence. The DNA sequence feature is often computed by combining sequence-based DNA embedding with various other features and the clustering algorithms are often complicated and strongly correlated with the features they utilize.
In our evaluation, to create a fair environment for DNA embedding benchmarking, instead of relying on any existing tool, we implement the modified K-medoid clustering algorithm proposed in Kang et al. [2015] for metagenomics binning due to its simplicity and effectiveness. Algorithm 1 describes the unsupervised clustering algorithm we used for metagenomics binning, where $\texttt{s}({E_{i},E_{j}})$ represents the cosine similarity of two vectors $E_{i}$ and $E_{j}$. Selection of threshold $\gamma$. As shown in Algorithm 1, the threshold $\gamma$ is the most important hyperparameter that greatly impacts the final binning results. A high threshold results in small and dense clusters while a low threshold results in large yet sparse clusters. Since different models generate embeddings with distinct distributions, a fixed threshold (e.g., $0.9$) could be too high for one model yet too low for another one. In practice, massive hyperparameter searches are needed to determine the best threshold for each model on different datasets. Due to the large size of our experiments and the various types of models we used, an automatic way is needed to fairly choose the threshold for each model on each dataset. For each metagenomics binning dataset, we use the dataset from the same source (e.g., Marine) as it with ID 0 to compute a threshold for each model on it, Specifically, we generate embeddings for each DNA sequence in the dataset and compute the similarities between each DNA sequence and its species center (i.e., the average of all the DNA sequence belongs to this species). The $70$ percentile of all the similarities is used as the threshold. Other hyperparameters. We set minimum bin size $m=10$, number of steps $Z=1000$, and number of iterations $T=3$. We also experimented with $T=3,4,5$ and $60,70,80,90$ percentile of all the similarities is used as the threshold $\gamma$, and found that the results are robust to these hyperparameters.
Algorithm 1 Modified $K$-Medoid Clustering
[…]
3:for step $z=1$ to $Z$ do
4:  Select seed index $s=\arg\max_{s’}d_{s’}$ and corresponding seed $E_{s}$
5:  for iteration $t=1$ to $T$ do
[…]
8:  end for
9:  Set $p_{i}\leftarrow z,d_{i}\leftarrow 0$ for each $i\in\mathcal{I}$
[…]
11:end for
12:for step $z=1$ to $Z$ do
13:  Find indices $\mathcal{I}$ where $p_{i}=z$ for each $i\in\mathcal{I}$
14:  if $|\mathcal{I}|<m$ then
15:   Set $p_{i}\leftarrow-1$ for each $i\in\mathcal{I}$
16:  end if
17:end forReturn: predictions $\bm{p}$
Following Kang et al. [2015, 2019], we formulate this problem as identifying non-overlapping clusters of DNA sequences from the entire dataset, where each cluster of sequence is considered as an identified species. We then compare the predicted clusters with the true labels to count the number of species that have been successfully identified. A species is considered to be successfully identified if the F1 score of this species is over $0.5$. We compare different models by the number of species they identify with different levels of F1 scores (e.g., 0.50.60.5-0.6, 0.80.90.8-0.9). We only use the DNA embeddings as the feature of each DNA sequence.
Figure 4 shows the models’ performance on $6$ metagenomics binning datasets. As shown in the figure, similar to our observation in clustering, DNABERT-S identifies twice the number of species with an F1 score of over $0.5$ compared to the strongest baseline, showing its great capability in tackling important real-world biology challenges. Notably, DNABERT-S identifies a large number of species with an F1 score over $0.9$. indicating its capability to accurately segregate different species in the embedding space, aligning with our observation in Figure 1. In the Synthetic datasets, where the sequences are error-less (extracted from reference genome) and the number of sequences in each species is more balanced, DNABERT-S recovers over $80\%$ of the species with an F1 score of over $0.5$ purely based on the DNA sequences themselves. In more realistic datasets such as Marine and Plant, where noise (e.g., error from sequences) exists in DNA sequence and species size is highly imbalanced, DNABERT-S is still able to recover about $40\%$ of the species with an F1 score of over $0.5$.

### v3

@@ 3.2 Curriculum Contrastive Learning ($\text{C}^{2}$)
^$\hat{\ell}(f_{m}(h_{i}^{m}),v_{i}^{mix})=-\sum_{n=1}^{B}v_{i,n}^{mix}\log\frac{\exp\left(\operatorname{s}\left(f_{m}(h_{i}^{m}),f(x_{n^{+}})\right)/\tau\right)}{\sum_{j=1}^{B}\alpha_{ij^{+}}\exp\left(\operatorname{s}\left(f_{m}(h_{i}^{m}),f(x_{j^{+}})\right)/\tau\right)},$
@@
$\alpha_{ij^{+}}=\frac{\exp\left(\operatorname{s}\left(f_{m}(h_{i}^{m}),f(x_{j^{+}})\right)/\tau\right)}{\frac{1}{B-1}\sum_{k=1,k\neq i}^{B}\exp\left(\operatorname{s}\left(f_{m}(h_{i}^{m}),f(x_{k^{+}})\right)/\tau\right)}.$
@@ 3.3 Implementation
In the $\text{C}^{2}$ method, we set temperature $\tau$ as $0.05$ and hyperparameter $\alpha$ as $1.0$. We train the model for one epoch in phase I using loss Eq. (2) and for two epochs in phase II using loss Eq. (3). We use mean pooling of the last hidden states of all the tokens as the DNA embedding. We employ the Adam optimizer [Kingma and Ba, 2014], with a learning rate of $63e-$ 6 and batch size of $48$. We save the model every 10000 training steps and select the best one based on the validation loss in the validation dataset. We use the pre-trained DNABERT-2 [Zhou et al., 2023] as the starting point of contrastive training. We also conduct parallel experiments with HyenaDNA [Nguyen et al., 2023]. In Section 5.6, we show that DNABERT-2 outperforms HyenaDNA after the same contrastive training. The training of DNABERT-S takes approximately 48 hours on 8 NVIDIA A100 80GB GPUs.
@@ 4 Data
Training
Each training sample of DNABERT-S is a pair of non-overlapping DNA sequences extracted from the same species. We focus on microbe species since the genetic diversity within this group provides a rich substrate for examining the nuances of species differentiation. The dataset is constructed with the reference genomes from GenBank [Benson et al., 2012]. We obtained $47923$ pairs from $17636$ viral genomes, $1$ million pairs from $5011$ fungi genomes, and $1$ million pairs from $6402$ bacteria genomes. We randomly selected $2$ million pairs from the entire $2047923$ pairs of DNA sequences to construct the training data. The rest pairs are treated as validation data. All the DNA sequences are 10000 bp in length.
Evaluation
Our evaluation spans $14$ long-read datasets from the Critical Assessment of Metagenome Interpretation (CAMI) II [Meyer et al., 2022] challenge benchmark and $9$ synthetic datasets from reference genomes. CAMI2 is one of the most comprehensive and rigorous benchmarks for metagenomics research. The datasets in CAMI2 are designed to mimic realistic microbiome environments and include a vast array of both new and known genomes, as well as plasmids and viruses. It aligns our study with real-world ecological and biological scenarios, providing a robust and contextually relevant evaluation for the DNA embedding models. We utilize $7$ datasets of long-read contigs respectively from the Marine and Plant-associated environments, where each dataset consists of $150$k-$200$k DNA sequences belonging to about $100-750$ different species sampled from $1680$ microbial genomes and $599$ circular elements. We also create $9$ Synthetic datasets by randomly extracting DNA sequences from fungi and viral reference genomes that do not overlap with our training data. Table 5 in Appendix 5 shows the statistics of the datasets we used for evaluation.
@@ 5 Experiments
We evaluate the model in a series of tasks, including: 1) metagenomics binning that identifies species from a mixture of sequences from an unknown number of species; 2) species clustering given the number of species; 3) species classification with a few labeled samples; and 4) long-read classification given reference genomes. Since the CAMI2 datasets are highly imbalanced, for the clustering and classification tasks, we filtered the datasets to eliminate species with fewer than $100$ sequences and only kept $100$ sequences for each species, resulting in a set of perfectly balanced datasets. For the metagenomics binning problem, to mimic real-world scenarios, we do not balance the data. Instead, following Kang et al. [2015], we only keep DNA sequences longer than $2500$bp and filter out species with fewer than $10$ sequences.
In this section, we present experimental design and empirical results. We introduce baselines in Sec. 5.1 and respectively present the results of clustering in Sec. 5.2, metagenomics binning in Sec. 5.3, and classification in Sec. 5.4. In Sec. 5.5, we present ablation studies on C2LR and the proposed Manifold Instance Mixup training objective. Comparison with an alignment-based method for species classification is presented in Appendix 6. We also provide empirical analysis on varying backbone models, different input lengths, reduced feature dimensions, scenarios with abundant training data, various other types of tasks, and results with error bars. Due to space limitations, we present them in Appendix C. For all tasks involving randomness, we perform $5$ independent runs with different random seeds for each model and report the averaged results.
@@ 5.1 Baselines
We compare our model with four lines of work to examine its effectiveness. TNF, TNF-K, and TNF-VAE are the most widely used DNA embedding methods in metagenomics binning tools [Kang et al., 2015, 2019; Nissen et al., 2021]. TNF represents Tetra-Nucleotide Frequency, which uses the appearance frequency of each unique 4-mer ($2564^{4}=$ 256 in total) in a DNA sequence as its embedding. TNF-K [Nissen et al., 2021] reduces TNF to $103$-dimension with a linear kernel, which utilizes DNA characteristics to reduce the correlations among different dimensions of the original TNF feature. TNF-VAE trains a Variational Autoencoder [Kingma and Welling, 2013] using TNF as input to extract features. DNA2Vec [Ng, 2017] learns pre-trained K-mer embedding. We set $K=4$ to make it directly comparable with TNF and use the average of the $4$-mer embeddings as the DNA embedding. DNABERT-2 [Zhou et al., 2023], HyenaDNA [Nguyen et al., 2023], and NT-v2 (Nucleotide Transformer-v2) [Dalla-Torre et al., 2023] are representative genome foundation models. We use the average of the last hidden states as the DNA embedding. For evaluations, we utilized the respective pre-trained models from Huggingface ModelHub, specifically zhihan1996/DNABERT-2-117M, LongSafari/hyenadna-medium-450k-seqlen-hf, and InstaDeepAI/nucleotide-transformer-v2-100m-multi-species. DNA-Mutate, DNA-Dropout, and DNA-Double are variants of DNABERT-S, with the same hyperparameters but different positive pair construction strategies in contrastive training. DNA-Mutate views the same DNA sequence before and after random mutation (i.e., swap and delete $5\%$ of nucleotides) as a positive pair. DNA-Dropout passes the same DNA sequence through the embedding model (with a dropout rate $0.1$) twice and views the two distinct embeddings as a positive pair. DNA-Double views a DNA sequence and its reverse complementary (e.g., AATTC v.s. TTAAG) as a positive pair. We also compare the proposed curriculum contrastive learning framework and MI-Mix training objective with well-established contrastive learning methods, including Weight SimCLR [Zhang et al., 2021], i-Mix [Lee et al., 2020] and Supervised Contrastive Learning (SupCon) [Khosla et al., 2021]. Results on different training objectives are presented in Section 5.5.
@@ 5.2 Clustering
In this task, we evaluate the embedding quality by how well a standard clustering algorithm can distinguish and cluster different species based on the embedding. To reduce the effects of other factors, we assume the number of species is known in this task. For each dataset, we compute the embedding of each DNA sequence and perform K-means clustering by setting the num_clusters as the number of species that exist in this dataset. We employ the Adjusted Rand Index (ARI) as the evaluation metric. ARI is a measure of the similarity between two data clusterings, adjusted for chance, providing a normalized index that ranges from $-1$ to $1$; the higher, the better.
Table 1 shows the models’ performance on clustering. As shown in the table, DNABERT-S consistently achieves the best performance on all the datasets and doubles the performance of the strongest existing method on average. Among all the baselines, TNF and its variant TNF-K achieve the best performance, explaining their wide usage in metagenomics binning. Yet, TNF’s performance is heavily limited since it is not learnable. TNF-VAE represents a naive algorithm that enables learning with TNF, yet it leads to big performance degradation, potentially resulting from the large gap between its training objective and the specific downstream application. Similarly, pre-trained Kmer embeddings from DNA2Vec also fail to effectively cluster different species.
Existing genome foundation models training with language modeling objectives, such as HyenaDNA and DNABERT-2, despite their good performance on labeled datasets, also fail to generate representative embedding without fine-tuning. The phenomenon that pre-trained foundation models underperform descriptive textual features in generating embedding for clustering and retrieval is also observed in the field of natural language processing [Reimers and Gurevych, 2019].
@@
5.3 Metagenomics Binning
Metagenomics binning is a crucial process in microbial ecology, involving the categorization of DNA sequences into groups that represent individual species. State-of-the-art metagenomics binning method [Kang et al., 2015, 2019; Nissen et al., 2021] always formulate this problem as a clustering problem with an unknown number of clusters based on the feature of each DNA sequence. The DNA sequence feature is computed by combining sequence-based DNA embedding with various other features and the clustering algorithms are often complicated and strongly correlated with the features they utilize. In our evaluation, to create a fair environment for DNA embedding benchmarking, instead of relying on any existing tool, we implement the modified K-medoid clustering algorithm proposed in Kang et al. [2015] for metagenomics binning due to its simplicity and effectiveness. Algorithm 1 describes the metagenomics binning algorithm. Following Kang et al. [2015, 2019], we formulate this problem as identifying non-overlapping clusters of DNA sequences from the entire dataset, where each cluster of sequence is considered as an identified species. We iteratively identify the densest point in the embedding space and take all the sequences that are close (determined by a learned threshold) to it as the group of the sequences that belong to the same species. The embeddings of the taken sequences are removed from the embedding space. The iteration ends as there are no regions that contain enough number sequences within the threshold. This algorithm, evaluates how well the embedding method clusters and aggregates different species within the embedding space. We then compare the predicted clusters with the true labels to count the number of species that have been successfully identified. A species is considered to be successfully identified if the F1 score of this species is over $0.5$. We compare different models by the number of species they identify with different levels of F1 scores (e.g., $0.5-0.6$, $0.8-0.9$). We only use the DNA embeddings as the feature of each DNA sequence.
Figure 3 shows the models’ performance on $6$ metagenomics binning datasets. As shown in the figure, similar to our observation in clustering, DNABERT-S identifies twice the number of species with an F1 score of over $0.5$ compared to the strongest baseline, showing its great capability in tackling important real-world biology challenges. Notably, DNABERT-S identifies a large number of species with an F1 score over $0.9$. indicating its capability to accurately segregate different species in the embedding space, aligning with our observation in Figure 1. In the Synthetic datasets, where the sequences are error-less (extracted from reference genome) and the number of sequences in each species is more balanced, DNABERT-S recovers over $80\%$ of the species with an F1 score of over $0.5$ purely based on the DNA sequences themselves. In more realistic datasets such as Marine and Plant, where noise (e.g., error from sequences) exists in DNA sequence and species size is highly imbalanced, DNABERT-S is still able to recover $40\%$ of the species with an F1 score of over $0.5$.
@@ 5.4 Classification
In this task, we evaluate the embedding quality by how well a simple model can classify different species based on a few labeled embeddings. We conduct experiments with a linear regression model and a non-linear multi-layer perceptron (MLP) to examine the embeddings’ linear and non-linear descriptiveness. We present results on linear classification in this section. The results in non-linear settings are consistent with those in linear ones. Due to space limits, we present them in Table 6 and 7 in the Appendix. As shown in Table 5, all the datasets we use for classification consist of $100$ DNA sequence for each species. We first compute the embedding of each DNA sequence with each model. In each evaluation run, we independently select $80$ embeddings from each species to form the test set. For the rest DNA sequences, we respectively sample $1$, $2$, $5$, $10$, and $20$ embeddings from each species to form the training set. A Logistic Regression model is trained on the training set and evaluated on the test set. We use the macro F1 score as the evaluation metric.
Figure 4 shows the models’ performance on $6$ datasets. The remaining results are consistent and are presented in Appendix C.3. As shown in the figure, DNABERT-S consistently achieves the best performance. DNABERT-S achieves better performance than the strongest baseline with only $20\%$ of training data. For example, with only $2$ training samples per category, DNABERT-S achieves higher F1 scores than the strongest baseline with $10$ training samples. With the same amount of training samples, DNABERT-S outperforms the baselines by a large gap. Notably, in the Synthetic datasets, where none of the species are seen during the contrastive training, a linear model trained with DNABERT-S embeddings achieves an F1 score of over $0.8$ in 200 classes classification with only $5$ labeled samples in each species, showing DNABERT-S’s capability in generalizing well on unseen data. To further validate the generalizability of DNABERT-S on more distinct datasets, we compile three datasets, including genomes from invertebrate, protozoa, and mammalian species, that are largely different from the training species (microbial genomes) of DNABERT-S. As illustrated in Section C.5, DNABERT-S also achieves good performance in these datasets.